# supplement

### v1

@@ 6.3 Comparison of point forecast accuracy
In Table LABEL:tab:1, averaged across 47 prefectures, we compare the point forecast accuracy among the functional time-series models. From the summary statistics of the KLD and JSD, the HDFPCA model offers the most accurate point forecasts for the female data. For the male data, the MFTS and MLFTS models provide the most accurate point forecasts. For the UFTS, MFTS, MLFTS, and FANOVA models, the number of retained functional principal components was determined by the EVR criterion. For the HDFPCA model, we follow the default setting of Gao et al. (2019), where the number of retained components is six and two in the first and second stages, respectively. In Appendix A, we also provide the results for $K=6$.
Table 1: Averaged across 47 prefectures, we evaluate and compare the point forecast accuracy, measured by the KLD and JSD, among the functional time-series models. The number of components is determined by the EVR criterion.
@@
KLD 1 0.006 0.010 0.006 0.006 0.008 0.005 0.005 0.004 0.005 0.004
2 0.007 0.011 0.006 0.007 0.010 0.006 0.006 0.005 0.005 0.005
3 0.009 0.013 0.007 0.008 0.010 0.007 0.007 0.005 0.006 0.006
4 0.011 0.014 0.009 0.010 0.011 0.007 0.007 0.006 0.007 0.006
5 0.013 0.016 0.010 0.012 0.012 0.008 0.008 0.007 0.008 0.007
6 0.016 0.017 0.011 0.013 0.010 0.009 0.009 0.008 0.008 0.008
7 0.019 0.019 0.013 0.016 0.015 0.011 0.010 0.008 0.009 0.009
8 0.024 0.021 0.016 0.019 0.012 0.012 0.011 0.010 0.010 0.010
9 0.030 0.024 0.019 0.024 0.018 0.013 0.013 0.011 0.011 0.011
10 0.037 0.028 0.023 0.029 0.011 0.014 0.014 0.012 0.012 0.012
11 0.042 0.033 0.025 0.034 0.024 0.015 0.016 0.011 0.012 0.014
12 0.049 0.027 0.029 0.037 0.009 0.016 0.011 0.011 0.011 0.016
13 0.058 0.032 0.037 0.046 0.031 0.017 0.012 0.014 0.013 0.019
14 0.072 0.039 0.045 0.055 0.011 0.018 0.014 0.016 0.015 0.023
15 0.098 0.047 0.061 0.076 0.045 0.019 0.015 0.021 0.021 0.032
16 0.130 0.059 0.091 0.104 0.015 0.020 0.017 0.031 0.028 0.044
Mean 0.039 0.026 0.026 0.031 0.016 0.012 0.011 0.011 0.011 0.014
Median 0.027 0.023 0.017 0.021 0.011 0.012 0.011 0.010 0.011 0.010
JSD 1 0.038 0.048 0.037 0.044 0.038 0.035 0.034 0.031 0.033 0.034
2 0.042 0.051 0.040 0.046 0.041 0.038 0.036 0.034 0.035 0.036
3 0.045 0.054 0.043 0.049 0.044 0.040 0.039 0.036 0.037 0.038
4 0.051 0.057 0.046 0.048 0.048 0.043 0.041 0.038 0.040 0.040
5 0.056 0.060 0.050 0.056 0.052 0.046 0.043 0.041 0.042 0.042
6 0.061 0.061 0.053 0.050 0.056 0.049 0.045 0.043 0.044 0.043
7 0.066 0.063 0.057 0.062 0.060 0.052 0.047 0.045 0.047 0.045
8 0.073 0.066 0.061 0.053 0.065 0.055 0.049 0.047 0.049 0.047
9 0.081 0.070 0.067 0.070 0.071 0.057 0.052 0.049 0.052 0.049
10 0.089 0.074 0.073 0.054 0.078 0.060 0.054 0.052 0.056 0.051
11 0.099 0.079 0.078 0.080 0.086 0.062 0.056 0.053 0.061 0.052
12 0.109 0.080 0.085 0.053 0.095 0.063 0.054 0.055 0.066 0.053
13 0.121 0.086 0.095 0.093 0.105 0.065 0.057 0.059 0.073 0.057
14 0.134 0.092 0.104 0.056 0.114 0.068 0.060 0.063 0.079 0.060
15 0.159 0.100 0.123 0.112 0.136 0.071 0.062 0.072 0.094 0.070
16 0.188 0.115 0.151 0.066 0.162 0.073 0.065 0.086 0.113 0.083
Mean 0.088 0.072 0.073 0.062 0.078 0.055 0.050 0.050 0.058 0.050
Median 0.077 0.068 0.064 0.055 0.068 0.056 0.050 0.048 0.051 0.050
@@
Figure 5 presents two heatmaps comparing seven functional time-series models across 16 forecast horizons for both female and male data. For each horizon, we count how often each method produces the smallest point forecast errors, as measured by the KLD. Choosing $K=6$ has a slight advantage compared to the one based on the EVR criterion. For shorter forecast horizons, the UFTS, MFTS, and MLFTS models perform equally well for the female data. However, for relatively longer forecast horizons, the HDFPCA model is recommended. In the case of male data, the MLFTS model is generally preferred; only at the shorter forecast horizon is $K=6$ advantageous. For relatively longer forecast horizons, the FANOVA with the EVR criterion is recommended.
@@
Figure 5: Heatmaps comparing the point forecast accuracy, measured by the KLD, between nine functional time-series models across 16 forecast horizons for the testing female and male data. Since there are 47 prefectures, each row sums to 47. A method appearing more than five times is highlighted in green, while those appearing fewer than five times are shown in red. If a method appears exactly five times, it is displayed in white.
@@ 6.4 Interval forecast error metrics
To evaluate interval forecast accuracy, we consider the CPD between the empirical coverage probability (ECP) and nominal coverage probability, as well as the mean interval score of Gneiting & Raftery (2007). For each year in the forecasting period, the $h$-step-ahead prediction intervals are computed at the $100(1-\alpha)\%$ nominal coverage probability. We consider the common case of the symmetric $100(1-\alpha)\%$ prediction intervals, with lower and upper bounds that are quantiles at $\alpha/2$ and $21-\alpha/$ 2, denoted by $\widehat{d}_{n+\xi,s}^{g,\text{lb}}(u)$ and $\widehat{d}_{n+\xi,s}^{g,\text{ub}}(u)$. The ECP and CPD are defined as
$\text{ECP}_{h}$ $=\frac{1}{111\times(16-h)}\times\sum_{\xi=h}^{15}\sum_{u=1}^{111}\mathds{1}\left\{\widehat{d}_{\nu+\xi,s}^{g,\text{lb}}(u)\leq d_{\nu+\xi,s}^{g}(u)\leq\widehat{d}_{\nu+\xi,s}^{g,\text{ub}}(u)\right\},$ (10)
$\text{CPD}_{h}$ $=\frac{1}{111\times(16-h)}\times\sum_{\xi=h}^{15}\sum_{u=1}^{111}\left[\mathds{1}\{d_{\nu+\xi,s}^{g}(u)>\widehat{d}_{\nu+\xi,s}^{g,\text{ub}}(u)\}+\mathds{1}\{d_{\nu+\xi,s}^{g}(u)<\widehat{d}_{\nu+\xi,s}^{g,\text{lb}}(u)\}\right],$
where $h=1,\dots,15$, $v$ denotes the end of training set when we calibrate the CPD or interval score, and $\nu$ denotes the end of validation set when we evaluate the interval forecast accuracy.
@@
$\overline{\text{CPD}}$ $=\frac{1}{15}\sum^{15}_{h=1}\text{CPD}_{h},$
@@
$\overline{S}_{\alpha}(h)=\frac{1}{111\times(16-h)}\times\sum^{15}_{\xi=h}\sum^{111}_{u=1}S_{\alpha,\xi}\left[\widehat{d}_{\nu+\xi,s}^{g,\text{lb}}(u),\widehat{d}_{\nu+\xi,s}^{g,\text{ub}}(u),d_{\nu+\xi,s}^{g}(u)\right].$
@@
$\overline{S}_{\alpha}$ $=\frac{1}{15}\sum^{15}_{h=1}\overline{S}_{\alpha}(h),$
$\text{M}[S_{\alpha}]$ $=\text{median}[\overline{S}_{\alpha}(h)].$
@@ 6.5 Comparison of interval forecast accuracy
In Table LABEL:tab:2, averaged across 47 prefectures, we compare the interval forecast accuracy among the functional time-series models using the EVR criterion. From the summary statistics of the CPD and interval score at the level of significance $\alpha=0.2$, the MFTS offers the smallest CPD value for the female data, while the UFTS and MLFTS offer the smallest CPD values for the male data. In terms of the interval scores, the HDFPCA provides the smallest interval scores for the female data, while the MLFTS gives the smallest interval scores for the male data. In Appendix B, we also provide the results for $K=6$.
Table 2: Averaged across 47 prefectures, we evaluate and compare the interval forecast accuracy, as measured by the CPD and interval score, among the seven functional time-series models at the level of significance $\alpha=0.2$. The number of components is determined by the EVR criterion.
@@
$\overline{\text{CPD}}_{0.2}$ 1 0.061 0.065 0.043 0.075 0.072 0.056 0.085 0.044 0.074 0.059
2 0.067 0.058 0.045 0.088 0.053 0.058 0.088 0.047 0.079 0.071
3 0.071 0.054 0.052 0.097 0.074 0.055 0.090 0.052 0.088 0.073
4 0.081 0.051 0.058 0.117 0.076 0.054 0.094 0.057 0.093 0.084
5 0.087 0.048 0.065 0.122 0.091 0.056 0.096 0.061 0.098 0.086
6 0.086 0.045 0.069 0.128 0.100 0.058 0.095 0.058 0.095 0.090
7 0.090 0.047 0.071 0.130 0.069 0.057 0.091 0.056 0.087 0.084
8 0.095 0.046 0.074 0.131 0.107 0.060 0.085 0.056 0.089 0.075
9 0.096 0.050 0.080 0.131 0.060 0.072 0.083 0.060 0.087 0.076
10 0.104 0.052 0.088 0.143 0.138 0.073 0.082 0.063 0.087 0.086
11 0.103 0.051 0.082 0.144 0.048 0.075 0.086 0.074 0.084 0.084
12 0.105 0.059 0.087 0.136 0.134 0.069 0.083 0.067 0.080 0.077
13 0.106 0.061 0.093 0.137 0.048 0.069 0.082 0.080 0.080 0.080
14 0.116 0.067 0.094 0.145 0.096 0.065 0.079 0.076 0.089 0.077
15 0.122 0.069 0.093 0.137 0.086 0.057 0.067 0.071 0.070 0.085
Mean 0.093 0.055 0.073 0.118 0.083 0.062 0.086 0.062 0.085 0.079
Median 0.095 0.052 0.074 0.131 0.076 0.058 0.085 0.060 0.087 0.080
$\overline{S}_{0.2}$ 1 255 393 246 279 314 271 270 213 263 243
2 299 417 279 328 344 298 296 237 283 270
3 345 445 312 384 369 325 322 259 300 292
4 412 476 353 470 361 356 351 286 321 321
5 477 511 397 551 464 390 376 309 345 349
6 549 543 429 639 381 419 397 331 358 370
7 626 565 476 734 537 445 412 353 375 387
8 721 600 536 843 420 465 433 378 393 409
9 829 650 606 973 625 493 451 402 416 436
10 956 725 690 1133 476 525 481 432 439 488
11 1070 798 767 1271 793 543 506 430 455 519
12 1199 788 864 1429 500 570 467 451 460 560
13 1350 874 1000 1623 966 583 492 507 510 623
14 1501 945 1108 1823 546 661 547 582 571 669
15 1888 1122 1419 2224 1258 875 690 811 835 933
Mean 832 657 632 1000 557 481 433 399 420 458
Median 721 600 536 843 476 465 433 378 393 409
In Figure 6, we present two heatmaps comparing seven functional time-series models across 15 forecast horizons for both female and male testing data. For each horizon, we count how often each method produces the smallest interval forecast errors, as measured by the CPD and interval score. By examining the CPD results, we recommend the MFTS for the female data and the MLFTS model for the male data, especially for $K=6$. By examining the interval score results, we recommend the MFTS for shorter horizons and the HDFPCA for longer horizons in the female data. For the male data, the MLFTS generally provides the smallest interval scores.
@@
Figure 6: Heatmaps comparing the interval forecast accuracy at the level of significance $\alpha=0.2$, measured by the CPD and interval score, between seven functional time-series models across 15 forecast horizons for the female and male testing data. Since there are 47 prefectures, each row sums to 47. A method appearing more than five times is shown in green, while those appearing fewer than five times are shown in red. If a method appears exactly five times, it is shown in white.

### v2

@@ 6.3 Comparison of point forecast accuracy
In Table LABEL:tab:1, we compare the point forecast accuracy across the functional time-series models, averaged across 47 prefectures. Based on the summary statistics of the KLD and JSD, the HDFPCA model provides the most accurate point forecasts for the female data. This result indicates that the two-stage principal component analyses can adequately summarize the main features in the female data without much loss of information. The female data are relatively stable in comparison to the male counterpart. For the male data, the MFTS and MLFTS models provide the most accurate point forecasts. For the UFTS, MFTS, MLFTS, and FANOVA models, the number of retained functional principal components was determined by the EVR criterion. For the HDFPCA model, we follow the default setting of Gao et al. (2019), where the number of retained components is six and two in the first and second stages, respectively. In Appendix A, we also provide the results for $K=6$ as a sensitivity analysis.
Table 1: Averaged across 47 prefectures, we evaluate and compare the point forecast accuracy of the functional time-series models, measured by the KLD and JSD. The number of components is determined by the EVR criterion.
@@
KLD 1 0.004 0.010 0.004 0.006 0.007 0.004 0.004 0.003 0.005 0.003
2 0.007 0.013 0.006 0.007 0.007 0.005 0.006 0.005 0.005 0.003
3 0.008 0.015 0.007 0.009 0.009 0.006 0.007 0.005 0.006 0.004
4 0.010 0.017 0.009 0.010 0.006 0.006 0.007 0.006 0.007 0.005
5 0.013 0.019 0.010 0.012 0.013 0.007 0.008 0.007 0.007 0.006
6 0.015 0.021 0.011 0.013 0.007 0.008 0.009 0.007 0.008 0.007
7 0.018 0.023 0.012 0.015 0.019 0.009 0.010 0.008 0.008 0.009
8 0.022 0.025 0.014 0.018 0.008 0.010 0.011 0.009 0.009 0.012
9 0.027 0.027 0.017 0.022 0.025 0.011 0.012 0.010 0.010 0.015
10 0.033 0.030 0.021 0.026 0.010 0.012 0.014 0.011 0.011 0.018
11 0.040 0.034 0.024 0.033 0.032 0.012 0.015 0.012 0.012 0.022
12 0.048 0.039 0.030 0.040 0.014 0.013 0.017 0.014 0.013 0.026
13 0.059 0.047 0.036 0.047 0.039 0.014 0.020 0.015 0.014 0.030
14 0.068 0.040 0.042 0.061 0.019 0.014 0.012 0.015 0.017 0.035
15 0.089 0.054 0.060 0.080 0.044 0.015 0.015 0.021 0.022 0.039
16 0.118 0.070 0.079 0.106 0.019 0.015 0.019 0.029 0.030 0.040
17 0.154 0.084 0.093 0.120 0.038 0.015 0.020 0.034 0.034 0.039
Mean 0.043 0.033 0.028 0.037 0.019 0.010 0.012 0.012 0.013 0.019
Median 0.027 0.027 0.017 0.022 0.014 0.011 0.012 0.010 0.010 0.015
JSD 1 0.034 0.049 0.034 0.039 0.042 0.030 0.032 0.028 0.034 0.029
2 0.041 0.055 0.040 0.043 0.038 0.035 0.036 0.033 0.036 0.030
3 0.045 0.059 0.043 0.046 0.050 0.037 0.039 0.036 0.038 0.032
4 0.049 0.063 0.046 0.049 0.037 0.040 0.041 0.038 0.040 0.035
5 0.054 0.066 0.049 0.053 0.060 0.042 0.043 0.040 0.041 0.039
6 0.058 0.069 0.052 0.056 0.040 0.044 0.045 0.041 0.043 0.043
7 0.063 0.071 0.055 0.060 0.070 0.047 0.047 0.043 0.044 0.049
8 0.069 0.073 0.059 0.065 0.044 0.050 0.049 0.045 0.046 0.055
9 0.076 0.075 0.063 0.070 0.082 0.052 0.051 0.048 0.047 0.062
10 0.083 0.079 0.069 0.076 0.051 0.054 0.053 0.050 0.049 0.069
11 0.092 0.083 0.075 0.085 0.092 0.056 0.054 0.052 0.052 0.075
12 0.102 0.088 0.082 0.096 0.058 0.058 0.056 0.055 0.054 0.081
13 0.116 0.094 0.092 0.106 0.100 0.060 0.059 0.058 0.057 0.087
14 0.128 0.097 0.102 0.121 0.066 0.061 0.057 0.061 0.064 0.094
15 0.150 0.112 0.122 0.141 0.104 0.062 0.062 0.070 0.072 0.097
16 0.176 0.126 0.141 0.164 0.068 0.064 0.067 0.081 0.084 0.095
17 0.202 0.134 0.152 0.175 0.094 0.064 0.069 0.087 0.088 0.093
Mean 0.090 0.082 0.075 0.085 0.064 0.050 0.051 0.051 0.052 0.063
Median 0.076 0.075 0.063 0.070 0.060 0.052 0.051 0.048 0.047 0.062
@@
Figure 6 presents two heatmaps comparing seven functional time-series models across 16 forecast horizons for both female and male data. For each horizon, we count how often each method produces the smallest point forecast errors, as measured by the KLD. Choosing $K=6$ has a slight advantage compared to the one based on the EVR criterion. For shorter forecast horizons, the UFTS, MFTS, and MLFTS models perform equally well for the female data. However, for relatively longer forecast horizons, the HDFPCA model is recommended. For male data, the MLFTS model is generally preferred; only at the shorter forecast horizon is $K=6$ advantageous. For relatively longer forecast horizons, the FANOVA with the EVR criterion is recommended.
@@
Figure 6: Heatmaps comparing the point forecast accuracy, measured by the KLD, across 17 forecast horizons for nine functional time-series models on the testing female and male data. Since there are 47 prefectures, each row sums to 47. A method that appears more than 5 times is highlighted in green, while those that appear fewer than 5 times are shown in red. If a method appears exactly five times, it is displayed in white.
To formally assess whether the observed differences in point forecast accuracy obtained via the different presented methods are statistically significant, we further apply the Model Confidence Set (MCS) procedure (Hansen et al., 2011) based on the Kullback–Leibler divergence across all forecast horizons and prefectures. The detailed MCS results and corresponding heatmap visualizations are reported in Appendix D.
@@ 6.4 Interval forecast error metrics
To evaluate interval forecast accuracy, we consider the CPD between the empirical coverage probability (ECP) and nominal coverage probability, as well as the mean interval score of Gneiting & Raftery (2007). For each year in the forecasting period, the $h$-step-ahead prediction intervals are computed at the $100(1-\alpha)\%$ nominal coverage probability. We consider the common case of the symmetric $100(1-\alpha)\%$ prediction intervals, with lower and upper bounds that are quantiles at $\alpha/2$ and $1-\alpha/2$, denoted by $\widehat{d}_{n+\xi,s}^{g,\text{lb}}(u)$ and $\widehat{d}_{n+\xi,s}^{g,\text{ub}}(u)$. The ECP and CPD are defined as
$\text{ECP}_{h}$ $=\frac{1}{111\times(17-h)}\times\sum_{\xi=h}^{16}\sum_{u=1}^{111}\mathds{1}\left\{\widehat{d}_{\nu+\xi,s}^{g,\text{lb}}(u)\leq d_{\nu+\xi,s}^{g}(u)\leq\widehat{d}_{\nu+\xi,s}^{g,\text{ub}}(u)\right\},$ (10)
$\text{CPD}_{h}$ $=\frac{1}{111\times(17-h)}\times\sum_{\xi=h}^{16}\sum_{u=1}^{111}\left[\mathds{1}\{d_{\nu+\xi,s}^{g}(u)>\widehat{d}_{\nu+\xi,s}^{g,\text{ub}}(u)\}+\mathds{1}\{d_{\nu+\xi,s}^{g}(u)<\widehat{d}_{\nu+\xi,s}^{g,\text{lb}}(u)\}\right],$
where $h=1,\dots,16$, $\nu$ denotes the end of the training set when we calibrate the CPD or interval score, and $\nu$ denotes the end of the validation set when we evaluate the interval forecast accuracy.
@@
$\overline{\text{CPD}}$ $=\frac{1}{16}\sum^{16}_{h=1}\text{CPD}_{h},$
@@
$\overline{S}_{\alpha}(h)=\frac{1}{111\times(17-h)}\times\sum^{16}_{\xi=h}\sum^{111}_{u=1}S_{\alpha,\xi}\left[\widehat{d}_{\nu+\xi,s}^{g,\text{lb}}(u),\widehat{d}_{\nu+\xi,s}^{g,\text{ub}}(u),d_{\nu+\xi,s}^{g}(u)\right].$
@@
$\overline{S}_{\alpha}=\frac{1}{16}\sum^{16}_{h=1}\overline{S}_{\alpha}(h),\qquad\text{M}[S_{\alpha}]=\text{median}[\overline{S}_{\alpha}(h)].$
@@ 6.5 Comparison of interval forecast accuracy
In Table LABEL:tab:2, averaged across 47 prefectures, we compare the interval forecast accuracy among the functional time-series models using the EVR criterion. From the summary statistics of the CPD and interval score at the level of significance $\alpha=0.2$, the MFTS offers the smallest CPD value for the female data, while the UFTS and MLFTS offer the smallest CPD values for the male data. In terms of the interval scores, the UFTS and MFTS provide the smallest interval scores for the female data, while the FANOVA gives the smallest interval scores for the male data. In Appendix B, we also provide the results for $K=6$.
Table 2: Averaged across 47 prefectures, we evaluate and compare the interval forecast accuracy, as measured by the CPD and interval score, for the seven functional time-series models at the significance level $\alpha=0.2$. The number of components is determined by the EVR criterion.
@@
$\overline{\text{CPD}}_{0.2}$ 1 0.061 0.080 0.047 0.075 0.077 0.060 0.100 0.043 0.079 0.067
2 0.060 0.074 0.046 0.085 0.054 0.060 0.102 0.047 0.090 0.084
3 0.058 0.068 0.050 0.096 0.091 0.058 0.105 0.055 0.094 0.086
4 0.058 0.064 0.055 0.110 0.067 0.058 0.108 0.056 0.097 0.087
5 0.053 0.056 0.057 0.113 0.094 0.053 0.101 0.056 0.096 0.085
6 0.060 0.050 0.065 0.124 0.085 0.060 0.104 0.057 0.097 0.092
7 0.061 0.048 0.069 0.125 0.084 0.061 0.097 0.057 0.091 0.094
8 0.061 0.046 0.074 0.126 0.095 0.061 0.089 0.055 0.085 0.087
9 0.068 0.047 0.081 0.128 0.072 0.068 0.083 0.062 0.076 0.084
10 0.071 0.047 0.086 0.139 0.123 0.071 0.085 0.069 0.084 0.090
11 0.078 0.049 0.093 0.145 0.059 0.078 0.087 0.075 0.085 0.084
12 0.075 0.053 0.092 0.146 0.126 0.075 0.083 0.078 0.085 0.089
13 0.076 0.054 0.084 0.147 0.065 0.076 0.073 0.077 0.089 0.081
14 0.073 0.068 0.087 0.147 0.100 0.073 0.078 0.084 0.095 0.099
15 0.069 0.090 0.091 0.163 0.098 0.069 0.080 0.088 0.087 0.105
16 0.062 0.098 0.093 0.178 0.062 0.062 0.077 0.081 0.085 0.088
Mean 0.065 0.062 0.073 0.128 0.085 0.065 0.091 0.065 0.089 0.087
Median 0.062 0.055 0.078 0.127 0.084 0.062 0.088 0.060 0.088 0.086
$\overline{S}_{0.2}$ 1 268 256 433 287 325 256 279 218 258 248
2 295 302 466 336 335 295 306 243 279 281
3 317 347 494 389 402 317 332 265 293 306
4 345 397 524 450 357 345 359 285 312 327
5 368 445 551 508 457 368 379 300 321 344
6 398 510 591 587 394 398 401 319 338 370
7 420 583 618 672 547 420 415 339 354 394
8 443 675 651 780 450 443 432 361 371 420
9 465 763 688 885 633 465 446 386 386 441
10 488 875 750 1033 492 488 473 409 398 484
11 500 998 761 1191 775 500 491 439 421 513
12 520 1159 862 1377 542 520 518 461 448 585
13 527 1306 914 1572 1020 527 470 482 473 658
14 563 1533 1058 1877 617 563 521 571 535 783
15 627 1869 1225 2203 1432 627 588 700 621 924
16 843 2151 1456 2526 848 843 807 991 865 1139
Mean 481 481 784 1281 705 481 491 453 440 534
Median 443 447 688 885 471 443 446 439 421 431
In Figure 7, we present two heatmaps comparing seven functional time-series models across 16 forecast horizons for both female and male testing data. For each horizon, we count how often each method produces the smallest interval forecast errors, as measured by the CPD and interval score. Based on the CPD results, we recommend the MFTS for the female data and the MLFTS for the male data, especially for $K=6$. By examining the interval score results, we recommend the MFTS for shorter horizons and the HDFPCA for longer horizons in the female data. For the male data, the MFTS generally provides the smallest interval scores.
@@
Figure 7: Heatmaps comparing the interval forecast accuracy at the significance level $\alpha=0.2$, measured by the CPD and interval score, across seven functional time-series models and 16 forecast horizons for the female and male testing data. Since there are 47 prefectures, each row sums to 47. Methods that appear more than 5 times are shown in green, while those that appear fewer than 5 times are shown in red. If a method appears exactly five times, it is shown in white.
@@
6.6 Forecasting Age-Specific Survival Probabilities and Life Expectancy
The enhanced forecast accuracy of subnational life-table death counts ($\widehat{d}_{x}$) enables precise estimation of key demographic indicators, such as the age-specific probability of dying ($q_{x}$) and life expectancy ($e_{x}$). These metrics are derived by constructing a complete life table from the forecasted death counts ($d_{x}$). Specifically, we apply the standard life table relationships using the LifeTable and convertFx functions from the MortalityLaws package (Pascariu, 2025) in [Uncaptioned image]. The life table radix ($l_{0}$) is set to 100,000 for all forecasted tables. For illustration, Figure 8 shows the forecasted distributions of deaths, probabilities of dying, and life expectancy for the prefecture Hokkaido for the years 2007 to 2023, based on data from 1975 to 2006.
Figure 8: Forecasted life table metrics for Hokkaido, 2007–2023. Panels show age-specific deaths ($d_{x}$), probability of dying ($q_{x}$), and life expectancy ($e_{x}$) for females (F) and males (M).
6.7 Model Diagnostics
We illustrate model diagnostics using the FANOVA residuals to assess the weak functional white noise assumption, focusing on stationarity and conditional heteroscedasticity. Stationarity was tested using the $\mathbf{T}$-statistic-based method for functional time series (T_stationary in ftsa package in [Uncaptioned image], Hyndman & Shang, 2025), separately for female and male residuals (matrices $110\times 49$ per prefecture). Conditional heteroscedasticity was assessed using the functional conditional heteroscedasticity test (fCH_test in FTSgo package in [Uncaptioned image], Rice et al., 2020). Table 3 summarizes the $p$-values across the 47 prefectures. Since all $p$-values exceed 0.05, the residuals can be considered approximately stationary. Small $p$-values indicate some heteroscedasticity, but as our methods model the conditional mean, variance modeling is beyond the scope; functional GARCH-type models could be explored in future work.
Table 3: Summary of $p$-values from stationarity and conditional heteroscedasticity tests.
   Gender    Min    1st Qu.    Median    Mean    3rd Qu.    Max
Stationary test
   Female    0.060    0.078    0.085    0.0853    0.091    0.108
   Male    0.057    0.0705    0.079    0.0784    0.0855    0.099
Conditional heteroscedasticity test
   Female    0.0000    8.22e-15    1.64e-12    7.67e-04    2.58e-10    3.60e-02
   Male    3.85e-13    3.35e-09    6.56e-06    1.61e-02    2.13e-04    4.88e-01